OpenStudy (anonymous):

What is the maximum vertical distance between the line y = x + 72 and the parabola y = x2 for −8 ≤ x ≤ 9?What is the maximum vertical distance between the line y = x + 72 and the parabola y = x^2 for −8 ≤ x ≤ 9?

OpenStudy (anonymous):

distance definition: \[d =\sqrt{(x1-x2 )^{2}+(y1-y2)^{2}}\] \[y1=x+72\] \[y2=x^{2}\] \[d=x+72-x^{2}\] \[d(-8)=0\] \[d(-9)=18\] Maximum distance in [-8,-9] is 18 Attached is a graph